OpenStudy (anonymous):

f(x) = 1 - 6x. f(2x) can be expressed in the form a + bx. Find the values of a and b

OpenStudy (anonymous):

hi, what is this topic called

OpenStudy (anonymous):

Functions probably

OpenStudy (anonymous):

thanks i know this question was from oct, but are u doing it as well

OpenStudy (anonymous):

Nope... But to talk through it a bit: A function is where you input some value (eg, x, or 2x) and something is done to it: INPUT ----> [FUNCTION] -----> OUTPUT For the original function, f(x) = 1 - 6x: if we input 10, the function becomes: f(10) 1 - 60, so, the result is -59. So it follows that: f(2x) = 1 - 6(2x) = 1 - 12x a=1, b=12 hope this helps

OpenStudy (anonymous):

oops, a=1, b=-12